a car initially traveling at 35m/s takes 2.9 sec to readh an acceleration of 8.4 m/s2. how far does the car travel?
1
Expert's answer
2019-09-25T08:32:26-0400

The displacement of a car


"d=v_it+\frac{at^2}{2}"

"=35\:\rm m/s\times 2.9\: s+\frac{8.4\:\rm m/s^2\times ( 2.9\: s)^2}{2}"

"=137\:\rm m"